1.2x - y = 0
3x + 2y = 7

Answer by mananth(12270)

1.2 x -1 y = 0 .............1
3 x + 2 y = 7 .............2
Eliminate y
multiply (1)by 2
Multiply (2) by 1
2.4 x -2 y = 0
3 x + 2 y = 7
Add the two equations
5.4 x = 7
/ 5.4
x = 1.30 OR 9/7
plug value of x in (1)
1.2 x + -1 y = 0
1.56 + -1 y = 0
-1 y = 0 -1.56
-1 y = -1.56
y = 1.56 OR 14/9